A Level Further Mathematics 9231 Paper 42, May/June 2025

Question paper and mark scheme50 marks6 questions

Solve it on ExamPal

Practise

Untimed. Write on the paper and check as you go.

Start

Exam

Timed, with a clock you set, like sitting the real paper.

Start

Opens full screen.

Opening the question paper

Read as textThe questions and the mark scheme as plain text. Easier to read on a phone.

Question paper

9231/42 · May/June 2025

Download PDF
1

The manager of a hardware store is interested in whether there is a difference in the amount spent per customer on weekdays (x)comparedtoweekends(y). Random samples of 120 customers on weekdays and 80 customers on weekends are taken and the amount spent by each customer is recorded. The results are summarised as follows. ∑x=10 470∑(x−xˉ)2=12 283∑y=6560∑(y−yˉ)2=13 520 Test at the 1% significance level whether there is a difference in the mean amount spent per customer on weekdays compared to weekends. You should not assume that the population variances of the amounts spent on weekdays and weekends are equal. [7]

[7]
2

As shown in the diagram, the continuous random variable X has probability density function f given by f(x)={a0⩽x⩽5,b−cx5⩽x⩽8,0otherwise, where a, b and c are constants.

Figure 1 of 1 for question 2
a

Show that a=213 and find the values of b and c.

[3]
b

Find the mean of X.

[3]
c

Find the median of X.

[2]
d

The random variable Y is defined by Y=X2. Find the cumulative distribution function for Y.

[4]
3

Eggs in a supermarket are sold in boxes of six. A supermarket manager wishes to model the number of broken eggs in the boxes sold in the store. A random sample of 2000 boxes is taken and the number of broken eggs recorded. The observed frequencies are shown in the table below.

Number of broken eggs0123456
Observed frequency1844143110101
a

Use the data to estimate the probability that an egg is broken. Give your answer correct to 4 significant figures.

[1]
b

It is decided to carry out a goodness of fit test at the 0.5% significance level to determine whether a binomial distribution fits the data. The observed frequencies and the expected frequencies are given in the following table.

Number of broken eggs0123456
Observed frequency1844143110101
Expected frequency1831.3a6.0160.1190.0010.0000.000

Show that a=162.6 correct to 1 decimal place.

[1]
c

Carry out a goodness of fit test at the 0.5% level of significance to test whether a binomial distribution is a satisfactory model for the data.

[5]
d

Give a reason why a binomial distribution may not be a suitable model in this situation.

[1]
4

A researcher is interested in whether there is a difference between two schools in students’ aptitude for English. She randomly chooses 10 students from school X and 8 students of a similar age from school Y to take a written English test. The scores for the students from school X (x) and school Y (y) are summarised as follows. ∑x=612∑x2=40 104∑y=444∑y2=27 460 You should assume that the two distributions are normal and have the same population variance.

a

Find a 95% confidence interval for the difference in the mean scores for students from school X and students from school Y in the written English test.

[6]
b

Use the confidence interval you found in part (a) to explain why there is insufficient evidence at a 5% significance level to suggest that the English scores of students from school X and students from school Y are different.

[1]
5

Two jigsaw puzzles have the same number of pieces with identical shapes but have different pictures printed on them. One puzzle has a seaside picture and the other has a cartoon picture. A researcher believes that children will complete the cartoon puzzle more quickly. To test this belief, 10 children are randomly selected. The time taken in seconds for each child to complete each puzzle is recorded below.

ChildABCDEFGHIJ
Seaside182130193181192204184192180189
Cartoon161111195159202200168165145160
a

Carry out a Wilcoxon matched-pairs signed-rank test at the 5% significance level to test the researcher’s belief.

[6]
b

Show that using a paired-sample sign test at the 5% significance level would result in the opposite conclusion to that found in part (a).

[3]
c

It was later discovered that the experiment had been conducted such that each child completed the seaside puzzle first followed by the cartoon puzzle. Comment on the validity of using this experiment to test the researcher’s belief.

[1]
6

Y is a discrete random variable which takes the values 0, 2, 4, … . The probability generating function of Y is given by GY(t)=k1−at2.

a

Find k in terms of a.

[1]
b

Show that P(Y>2)=a2.

[3]
c

It is now given that a=0.2. Find the value of E(Y).

[2]

Mark scheme

9231/42 · May/June 2025

Download PDF

1 · 7 marks

  • B1

    H0:μx=μyH1:μx≠μy

  • M1

    sx2=103.218... sy2=171.139...

  • A1

    [sx2120+sy280=2.999]

  • M1

    Value of z obtained using correct form with their sx2 and their sy2

  • A1

    [z=]10470120−656080sx2120+sy280=3.03

  • M1

    '3.03' > 2.576, reject H0 (significant). OR P(Z>’3.03’)=0.00122<0.005, reject H0 (significant).

  • A1

    There is sufficient evidence to suggest the mean spend of customers is different on weekends compared to weekdays.

2a · 3 marks

  • B1

    5a+32a=1, leading to a=213.

  • M1

    Attempt to solve 2 independent equations in b and c or 3 independent equations in a, b and c.

  • A1

    b=1639,c=239

2b · 3 marks

  • M1

    [E(X)=]∫05213x dx+∫58x(1639−239x) dx

  • M1

    [113x2]05+[839x2−2117x3]58

  • A1

    4313=3.31

2c · 2 marks

  • M1

    213m=0.5 OR ∫0m213 dx=0.5

  • A1

    m=134=3.25

2d · 4 marks

  • B1

    213x,1639x−139x2(+k)

  • B1

    k=−2539

  • M1

    Both parts with y substituted (condone missing k), ignore limits.

  • A1

    G(y)={0y<0213y0⩽y⩽251639y−139y−253925<y⩽641y>64

3a · 1 mark

  • B1

    [0×1844+1×143+2×11+3×0+4×1+5×0+6×12000×6=]7480

3b · 1 mark

  • B1

    a=2000−(1831.3+6.016+0.119+0.001)=162.6 OR a=(61)×7480×(1−7480)5×2000=162.6

3c · 5 marks

  • B1

    H0:A binomial distribution is a satisfactory model for the data. H1:A binomial distribution is not a satisfactory model for the data.

  • M1

    (1844−1831.3)21831.3+(143−162.6)2162.6+(13−6.136)26.136[=0.08807+2.3626+7.6784] OR 184421831.3+1432162.6+1326.136−2000

  • A1

    10.1

  • M1

    [1 degree of freedom so critical value is 7.879.] '10.1' > 7.879, reject H0 (significant).

  • A1

    Sufficient evidence to suggest that the manager’s model is not satisfactory.

3d · 1 mark

  • B1

    Not independent or probability of “success” may not be constant or plausible reason in context (for example, higher number of boxes with more than 1 broken egg due to boxes being dropped).

4a · 6 marks

  • M1

    Σ(x−xˉ)2=40104−110×6122[=2649.6] Σ(y−yˉ)2=27460−18×4442[=2818]

  • A1

    sp2=2649.6+281810+8−2[=341.725]

  • B1

    2.120, SOI.

  • M1

    Correct form, must be a t value.

  • A1

    Correct un-simplified confidence interval, SOI.

  • A1

    [−12.9,24.3]

4b · 1 mark

  • B1

    Because zero lies inside the confidence interval.

5a · 6 marks

  • B1

    H0:ms=mc H1:ms>mc

  • M1

    Attempt at differences.

  • M1

    Attempt at (signed) ranks.

  • A1

    (P=51,Q=4)T=4

  • M1

    '4' ⩽10, reject H0 (significant).

  • A1

    Sufficient evidence to suggest that children, on average, complete the cartoon puzzle more quickly than the seaside puzzle. OR Sufficient evidence to support the researcher’s belief.

5b · 3 marks

  • B1

    [X∼B(10,0.5)]P(X⩽2)=0.0547

  • M1

    '0.0547' > 0.05, accept H0 (not significant).

  • A1

    In part 5(a) H0 is rejected, in part 5(b) H0 is accepted, so opposite conclusion reached.

5c · 1 mark

  • B1

    The order of the tests means that children may do the second test quicker (or slower) because, for example, they have had practice (or are tired).

6a · 1 mark

  • B1

    k=1−a

6b · 3 marks

  • M1

    (1−at2)−1=1+at2(+… ) OR P(Y=0)=GY(0)=k,P(Y=2)=12GY′′(0)=12(2ak)

  • 1 mark

    P(Y>2)=1−k(1+a) OR P(Y>2)=1−k−ak

  • A1

    P(Y>2)=1−(1−a)(1+a)=a2 OR P(Y>2)=1−(1−a)−a(1−a)=a2

6c · 2 marks

  • M1

    GY′(t)=0.32t(1−0.2t2)2

  • A1

    E(Y)=GY′(1)=0.32(1−0.2)2=0.5

Questions

  1. 1
    Question 17 marks
    Mark scheme ›
  2. 2
    Question 212 marks
    Mark scheme ›
  3. 3
    Question 38 marks
    Mark scheme ›
  4. 4
    Question 47 marks
    Mark scheme ›
  5. 5
    Question 510 marks
    Mark scheme ›
  6. 6
    Question 66 marks
    Mark scheme ›

More 9231 papers

Other papers from May/June 2025

Solve it on ExamPalStart